40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 345 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  247   Thu Jan 17 20:50:55 2008 tobinUpdateGeneralfiber coupling
Sam, John, and I matched the beam from an NPRO into a fiber on the SP table today. In doing so we used our GigE camera for a physics application for perhaps the first time, viewing the transmitted mode from the fiber during initial alignment. (I used my laptop running Windows and a 100 megabit switch.)
  246   Thu Jan 17 18:22:14 2008 AlbertoUpdateElectronicsRF Monitor Band-pass Filter
After we finalized the schematic for the RF monitor board based on buffered LC resonators, on Richard Abbott's suggestion to avoid the complication brought in by the fast op-amps, we gave another chance to the a passive configuration of the band-pass filter based on a Chebyshev topology. Rich and Ben gave me an old but very powerful software tool to design that kind of filters and showed me the way to circumvent many hassles in making RF test boards.

I made a test circuit for the 166MHz line (see attached schematic), using tunable inductors. The TF are also attached.
We get more than 20 dB of isolation after 33MHz (with a loss of only few dB at the resonance - it could be less), which is enough for all the other frequencies (33,133,199 MHz) but we would like more for the 166. We are going to add one or two extra orders to the filter.

We also have to understand the spike at about 320Mhz and eventually somehow get rid of it.


Alberto
Attachment 1: RF166Mhz.png
RF166Mhz.png
Attachment 2: Chebyshevb.png
Chebyshevb.png
Attachment 3: Chebyshev2b.png
Chebyshev2b.png
  245   Thu Jan 17 15:11:13 2008 josephbUpdateCamerasWorking on Malfalda
1) I can statically compile the ListCamera code (which basically just goes out and finds what cameras are connected to the network) on Malfalda and use that compiled code to run on Linux2 without a problem. Simply needed to add explicit links to libpthread.a and librt.a.
(i.e. -Bstatic -L /usr/lib/ -lpthread -Bstatic -L /usr/lib -lrt)

With appropriate static libraries, it should be possible to port this code to other linux machines even if we can't get it to compile on the target machine itself.

2)I've modified the Snap.cpp file so that it uses a packet size of 1000 or less. This simply involves setting the "PacketSize" attribute with the built in functions they provide in their library. After un-commenting some lines in that code, I was able to save tiff type images from the camera of up to 400x240 pixels on Malfalda. The claimed maximum resolution for the camera is 752x480, but it doesn't seem to work with the current setup. The max number of pixels seems to about 100 times the packet size. I.e. packet size of 1000 will allow up to 400x240 (96000) but not 500x240 (120,000). Not sure if this is an issue just with snap code or the general libraries used.

3)Will be working towards getting video running over the next day or so.
  244   Thu Jan 17 14:13:20 2008 robUpdateLSCWednesday's locking
Incremental progress on locking yet again. This time the handoff of DARM to the OMC worked, and progress halted at handing off control of the common mode to REFL166.
  243   Wed Jan 16 19:57:49 2008 tobinConfigurationPhotosISCT_EX
Here's a photo of the ISCT_EX table, for the purpose of planning our auxiliary laser arm locking scheme. Note the (undumped!) beam from the beamsplitter before QPDX (the leftmost gold-colored box); perhaps we could inject there.
Attachment 1: trx-annotated-small.jpg
trx-annotated-small.jpg
  242   Wed Jan 16 18:24:41 2008 ranaUpdateSUSETMY Watchdog
Because Steve keeps complaining about ETMY, I looked at some minute trend to see if there was something exotic happening at that time. It looks like there is some tremendous seismic activity to make it happen.

The trend shows that there is nothing special happening on the ETMX accelerometer or the ETMX suspension. At the same time, however, there is a huge jump in the ETMY sensors and therefore the watchdog signal. Whenever the watchdog value goes above 140, it trips.

After Andrey moves some accelerometers over to the Y end we can see the effect more directly.
Attachment 1: A.pdf
A.pdf
  241   Wed Jan 16 14:09:45 2008 robUpdateLSCtuesday's locking

I got a little further with the locking (PRFPMI) last night, after discovering that the cable going from the CM board to the MC board was unplugged at the MC side. This explains why we weren't able to engage the AO path last night. Tonight, I got up to the point where DARM is handed off to OMC transmission, a step which repeatedly failed.
Eventually I realized that although all the lights are the green, the OMC Trans signal was not being updated in the LSC's memory. I suspect this is because the c1ass machine was powered down. Work continues.
  240   Wed Jan 16 14:06:24 2008 robUpdateLSCmonday's locking
rob, tobin, johnnie

We did some locking work monday night, with decent progress. Working in the PRFPMI style, we managed to get through the part of the script that hands off the offset-CARM DOF to the MCL, but were not successful in engaging the AO path.

We also confirmed the problem with tdsread which prevent it from reading from multiple TLS (Three Lettered Subsystems) at the same time. Tobin traced this to a problem with the ezca library which tds uses, but it's not clear how to fix it. For now we just split the tdsread calls so that there are no multiple TLS calls. Tobin will report further on this.
  239   Tue Jan 15 13:15:27 2008 tobinUpdateEnvironmentlots of noise
They're throwing concrete around at the construction site.
  238   Mon Jan 14 23:11:26 2008 tobinConfigurationGeneralfiber
John and I removed the fiber that ran from the SP table to the cleanroom. We plan to build a MZ interferometer with this fiber inserted into one of the arms, for the purpose of measuring its phase noise.
  237   Mon Jan 14 14:41:09 2008 steveUpdateSUSetmy sus damping restored & mz relocked
Tree days trend of MZ HV drift is typical these days.
So as the etmy sus inability to hold damping for longer then 2-3 days.
Attachment 1: etmysus&mzhvtrend.jpg
etmysus&mzhvtrend.jpg
  236   Fri Jan 11 17:01:51 2008 pkpUpdateCamerasGigE again
So, here I detail all the efforts on the GigE so far

(1) The GiGE camera requires a minimum of 9 kb packet size, which is not available on mafalda or on my laptop ( both of which run Ubuntu and the Camera programs compile there). The programs which require smaller sizes work perfectly fine on these machines. I tried to statically compile the files on these machines so that I could then port them to the other machines. But that fails because the static libraries given by the company dont work.

(2) On Linux2, which lets me set a packet size as high as 9 Kb, it doesnt compile because of a GLIBC error. I tried updating the glibc and it tells me that the version already existing is the latest ( which it clearly is not). So I tried to uninstall GLIBC and reinstall it, but it wont let me uninstall (it == rpm) glibc, since there are a lot of dependencies. A dead end in essence.

Steps being taken

(1) Locally installing the whole library suites on linux2. Essentially install another version of gcc and g++ and see if that helps.
(2) IF this doesnt work, then the only course of action I can take is to cannibalize linux2's GigE card and put it on mafalda. ( I need permission for this Smile ).

Once again any suggestions welcome.
  235   Thu Jan 10 15:04:04 2008 steveUpdateSUSilluminator light effect on BS position
The bs chamber illuminator light was turned on this morning and left on.
Earlier on Rana noticed that the bs moved.
I follwed up to see what happened. I turned off oplev servo and tried to recenter on oplev pd
by adjusting pitch and yaw biases. It did not move. I looked at suspention and realized that the
illuminator was still on. I turned it off and to my amazement the the AP spot started flashing
Attachment 1: bssusilum.jpg
bssusilum.jpg
  234   Thu Jan 10 13:45:52 2008 PkpUpdateCamerasGLIBC Error
So, I have tried to compile the camera files which are in /cvs/cds/caltech/target/Prosilica/examples for the past 2 days now and have been unable to get rid of the following error. (specifically ListCameras.cpp, as it doesnt have any other libraries required, which unnecessarily complicates things)

../../bin-pc/x86/libPvAPI.so: undefined reference to `__stack_chk_fail@GLIBC_2.4'
collect2: ld returned 1 exit status
make: *** [sample] Error 1

I used to get this error on mafalda too, but I had fixed it by installing the latest version of the glibc libraries. Inspite of doing so on linux2, the error still persists. I suspect it had something to do with it being a FC3 machine. My own laptop, which also runs Ubuntu works fine too. The problem with these Ubuntu machines is that they dont let me set the packet sizes to 9 kb which is required by the camera. Linux2 does.

If anyone has any idea how to resolve this issue, please let me know.

Thanks
Pinkesh.
  233   Thu Jan 10 12:08:23 2008 steveUpdateSUSwhy did the BS move?
Attachment 1: bshopped.jpg
bshopped.jpg
  232   Thu Jan 10 10:38:02 2008 steveUpdateSUSetmy damping restored
The IST building onstruction has really started yesterday and continuing today with big heavy ground breaking
machinary. The MC is holding lock and the suspentions are hanging on.

ETMY does not like this.

SUS-MC2_LLPD_VAR monitor is a good indicator of seismic activity on this 12 days plot
Attachment 1: etmysus.jpg
etmysus.jpg
Attachment 2: sustrend16d.jpg
sustrend16d.jpg
  231   Thu Jan 10 00:12:01 2008 tobinSummaryLockingDR
[John, Tobin, Rana]

1. We found SUS_BS_SENSOR_UL to have a ratty signal and low DC value. Twiddling the cables at the BS satellite amplifier and vacuum feedthrough brought the signal back (to 0.667V), but it is still spiky, spiking up to a couple times per second. Rana suggested that these spikes might be scattered YAG laser light (as hypothesized in August). The spikes go away when we misalign the PRM or either ITM, and when we unlock the mode cleaner, lending credance to this theory. SUS_BS_SENSOR_UR also spikes, but much less frequently. We turned off C1:SUS-BS_ULSEN_SW2 and continued.

2. After dither alignment the oplev beams were centred and we were able to lock DRM plus either arm reliably (however locking in this state broke ./drstep_bang at the first ``Going DD''). We ran scripts/DRFPMI/bang/nospring/drdown_bang and were subsequently able to lock DRFPMI (i.e., full IFO) a couple times.

3. To do: Debug ./drstep_bang with just the DRM (no arms).
  230   Wed Jan 9 20:36:42 2008 GoUpdateTreasureMoney in lab
Go's Desk.
Attachment 1: DSC_0370.JPG
DSC_0370.JPG
  229   Wed Jan 9 20:29:47 2008 DmassAoGTMICoffee Carafe
If you have been using the coffee machine in the 40m, you may have noticed small brown flecks in your coffee mug. The carafe in the 40m has accumulated a layer of what is presumed to be old dried up coffee. When a small amount of water is swirled around in the bottom, flecks of the brown layer come off. Pictures below are of the inside of the carafe.

But does it provide adequate protection from 1064 light?
Attachment 1: DSC_0363.JPG
DSC_0363.JPG
Attachment 2: DSC_0365.JPG
DSC_0365.JPG
Attachment 3: DSC_0368.JPG
DSC_0368.JPG
  228   Wed Jan 9 10:47:15 2008 frickeUpdateComputersethernet wiring in office/control room
The electrical people have been here for three days, installing ethernet cables and drops in the 40m office area, in the old conduits where there was power and 10base2. Soon we will have reliable ethernet connections, instead of relying on switches hanging from the ceiling, etc!
  227   Tue Jan 8 15:20:17 2008 PkpUpdateCamerasGigE update
[Tobin , Pinkesh]

Finally we got the camera doing something (other than giving out its attributes). The only thing that seems to work so far is a program called AAviewer, which converts the image into an ASCII format and displays it on the screen. If you want to play around with it, log into mafalda (131.215.113.23) via rana.ligo.caltech.edu. Access /cvs/cds/caltech/target/Prosilica/bin-pc/x86/ and there should be a few programs in there, one of which is AAviewer, which requires you to get an IP address (which is 131.215.113.103) for the camera right now. (You can also get the IP information via the ListCameras program). The camera is physically in the 40m near the network rack.

Other programs dont seem to be working and its probably due to the network/packetsize issues. Since linux2 can change its packetsize to a higher number, I will get it to compile on linux2 for now and then give it a shot.
  226   Mon Jan 7 09:01:39 2008 steveUpdateSUSBS sus damping restored
The BS sus damping was lost at 8am Sunday morning.
Attachment 1: bssdl.jpg
bssdl.jpg
  225   Fri Jan 4 08:42:03 2008 steveUpdateSUSetmy trips again
ETMY sus damping tripped at 6am this morning
It was reset. We should put an accelerometer to the south end to see
the garbage dumping effect.
Attachment 1: etmy20m.jpg
etmy20m.jpg
Attachment 2: etmy120s.jpg
etmy120s.jpg
Attachment 3: etmysenV.jpg
etmysenV.jpg
  224   Thu Jan 3 12:38:49 2008 robBureaucracyTMISore throat

Quote:

I did not feel anything wrong yesterday, but unfortunately I have a very much sore throat today. I need to drink warm milk with honey and rinse my throat often today. So far I do not have other illness symptomes (no fever), so I hope that this small disease will not last for a long time, but I feel that it is better for me to cure my sore throat today at home (and probably it is safer for others in 40-m).

I took yesterday the book "Digital Signal Processing", so I have it for reading at home.

Hope to see you tomorrow.


I've added a new category--TMI--for entries along these lines.
  222   Thu Jan 3 09:55:11 2008 steveUpdateSUSetmy sus damping restored
ETMY watch dog was lost at midnight
Attachment 1: etmy12h.jpg
etmy12h.jpg
  221   Thu Jan 3 09:12:59 2008 steveUpdatePSLMZ servo
Here is MZ trend for one year and 40 days.
Now days it runs out of range on the low side.
This is the weakest link in the psl today.
Attachment 1: mz1y.jpg
mz1y.jpg
Attachment 2: mz40d.jpg
mz40d.jpg
  220   Thu Jan 3 08:53:55 2008 steveUpdateSUSetmy vs etmx
Rana have corrected sus gain damping setting of ETMY 8 days ago

gain settings: pos, pit, yaw & sd
etmx: 4,2,2,& -16
etmy: 4,2,2,& 50
Attachment 1: sus.jpg
sus.jpg
  219   Wed Jan 2 11:38:56 2008 tobinSummaryWIKI-40M Updatewiki spam
Our wiki keeps getting spammed. It turns out that you can upload attachments without a password, which is a bug. I'll ask David Barker whether there's a fix for this.
  218   Sun Dec 30 02:36:35 2007 pkpUpdateGeneralAnother update
So I followed suggestions 1 and 3 so far and have started writing up what all needs to be done in order to compile and use the camera. I wrote a program to ping the camera and get its properties and am working on a program to get an image. The reason why I want to write my own programs to do this, is that it will be easier to reuse and also to compile/use in the first place. The programs currently rest in /cvs/cds/caltech/target/Prosilica/ . Unfortunately I will be away for the next couple of days and will have another update on the 2nd.
  217   Thu Dec 27 18:18:56 2007 ranaUpdateComputersUpdate on GigE Camera

Quote:
So I finally got the linux software to compile on mafalda. I got the software to dump all the information regarding the camera onto a file. I tried to take a tiff snap and came up empty. So I looked at the configuration file and realized that the camera thinks that the frame-rate is a nan. Am reading up the manual to fix the frame-rate manually and then will attempt to take another snap.

All the files are in a folder called Prosilica in /home/controls/ on mafalda. All the executables are in /home/controls/Prosilica/bin-pc/x86/* . On another note, I am looking for a name for the camera. Any suggestions are welcome.


Suggestion #1: put this in the target area in a directory called /prosilica/. /home/controls is not backed up.

Suggestion #2: put a readme file in there on any work that was necessary to get it to compile.

Suggestion #3: make a wiki page for the camera with all the info that camera code developers will need
  216   Thu Dec 27 13:08:04 2007 ranaUpdateSUSETMY tripping
Here's a trend from the last 2 days of ETMX and ETMY. You can see that the damping gain increase
has made them now act much more alike. Problem fixed.
Attachment 1: Untitled.png
Untitled.png
  215   Thu Dec 27 12:12:02 2007 pkpUpdate Update on GigE Camera
So I finally got the linux software to compile on mafalda. I got the software to dump all the information regarding the camera onto a file. I tried to take a tiff snap and came up empty. So I looked at the configuration file and realized that the camera thinks that the frame-rate is a nan. Am reading up the manual to fix the frame-rate manually and then will attempt to take another snap.

All the files are in a folder called Prosilica in /home/controls/ on mafalda. All the executables are in /home/controls/Prosilica/bin-pc/x86/* . On another note, I am looking for a name for the camera. Any suggestions are welcome.
  214   Wed Dec 26 15:12:48 2007 ranaUpdateSUSETMY tripping
It turned out that the ETMY POS damping gain was set to 1.0 while the ETMX had 3.8.

I put both ETMs to a POS gain of 4 and then also set the PIT, YAW and SIDE gains for
ETMY. Let's see if its more stable now.

In the next week or so Andrey should have perfected his damping gain setting technique
and the numbers should be set more scientifically.
  213   Wed Dec 26 15:00:06 2007 ranaUpdateSUSETMY tripping
Steve mentioned to me that ETMY is still tripping more than ETMX. The attached DV plot
shows the trend of the watchdog sensors; essentially the RMS fluctuations of the shadow
sensors. (note** DV can make PNG format plots directly which are much better than JPG
when making plots and much smaller than PS or PDF when plotting lots of points).
Attachment 1: etm.png
etm.png
  212   Sat Dec 22 15:32:11 2007 tobinAoGEnvironmentants
Ants are everywhere: on the PSL table, on the circuit board I'm soldering...

I believe I have discovered their energy source.
Attachment 1: DSC_0361.JPG
DSC_0361.JPG
  211   Sat Dec 22 00:52:57 2007 tobinConfigurationPSLISS surgery
In an attempt to quell oscillations in the (unused) outer loop portion of the ISS, I shorted the "PD+" and "PD-" signals from the (nonexistent) outer-loop diodes, and soldered in 47pf compensation capacitors in C92 and C220. This seems to have eliminated oscillations seen at TP41 and TP42. There's still something amiss at TP30 and maybe TP20. Otherwise, the ISS seems happy. I can turn the gain slider to +15dB without saturation (with the HEPA off), though there seems to be less light on the diode (~3.9V) than a week or two ago.
  210   Fri Dec 21 20:32:25 2007 tobinUpdatePhotosGigE camera
I couldn't resist any longer: I plugged in the Prosilica GC 750 GigE camera and took it for a spin. This is the little CMOS camera which sends out video over gigabit ethernet.

There were no difficulties at all in getting it running. I just plugged in the power, plugged in ethernet, and put on a lens from Steve's collection. I downloaded the "Sample Viewer" from the Prosilica website and it worked immediately.

It turns out that "Kirk's" computer has not only a gigabit ethernet card, but a little gigabit ethernet switch. I plugged the camera into this switch. The frame rate is amazing. With the camera under fluorescent lights I thought I saw some wacky automatic gain control, but I think this ~10Hz flicker is aliasing of the 60 Hz room lighting.

I put the camera on the PSL table briefly and tried viewing the image from a laptop over the (54mbs) wireless network. This didn't work so well: you could get a couple frames out of the camera, but then the client software would complain that it had lost communications. It appeared that scattered 1064nm light did show up brightly on the camera image. There is a green ethernet cable currently stashed on the roof of the PSL that appears unused. We can try mounting the gigE CMOS cable in place of one of the CCD video cameras.

I did not try the Linux software.

The camera is currently set up at Kirk's desk, using the cool little tripod Rana got from CyberGuys.

This camera looks very promising! Also, in the test image attached below, a very unusual condition has been documented.
Attachment 1: robs_desk.png
robs_desk.png
  209   Thu Dec 20 21:58:28 2007 AndreySummaryComputer Scripts / ProgramsResults for 2 previous XARM measurements have been added

I attached results (plots) of yesterday's daytime and overnight measurements to the initial reports about those measurements.

These are ELOG entries # 201 and # 205.
  208   Thu Dec 20 21:57:34 2007 AndreyUpdateComputer Scripts / ProgramsMeasurements in XARM today

Today at 2PM I started a program, it should change the suspension gains in the interval from 1.0 to 3.8 with the step 0.2. Estimated running time is till 3.30AM coming night.

Results will be reported on Friday.

BELOW: ADDITION MADE ON FRIDAY EVENING.

Due to some unforeseen circumstances, I was unable to add results on Friday. I have so far accelerometer spectra only, which I add to this ELOG entry.

I have files with the measurement results, and I will process them after Christmas and add to this ELOG entry. I might not be in the lab on Dec. 24 and 25.
Attachment 1: Accelerom_ETMX.png
Accelerom_ETMX.png
Attachment 2: Accelerom_ITMX.png
Accelerom_ITMX.png
  207   Thu Dec 20 19:10:03 2007 waldmanUpdateOMCStressful reattachment of heater
Photos may follow eventually, but for now here's the rundown. I scraped the heater clean of the thermal epoxy using a clean razor blade. Then I stuffed a small piece of lint free cloth in the OTAS bore and wrapped the OMC in tin foil. With a vacuum sucking directly from the face of the OTAS, I gently scraped the glue off the OTAS aluminum. I wiped both the OTAS and the heater down with an isoproponal soaked lint-free cloth. I put a thin sheen of VacSeal on the face of the heater, wiping off the excess from the edges with a cloth. Then I clamped the heater to the OTAS using 2" c-clamps from the tombstone back to the heater front, making sure the alignment of the OTAS was correct (connector on the absolute bottom, concentric with the OTAS outer diameter). I added a second clamp, then beaded the outside of the joint with a little bit extra VacSeal, just for kicks. I'll leave it covered at least overnight, and maybe for a day or two.

sam
  206   Thu Dec 20 19:05:34 2007 waldmanHowToOMCHOWTO build front ends
For instance, to build the TPT front end code.

  • Save your file /cvs/cds/advLigo/src/epics/simLink/tpt.mdl
  • go to /cvs/cds/advLIGO on the TPT machine
  • do make clean-tpt tpt install-tpt
  • do rm /cvs/cds/caltech/chans/daq/C2TPT.ini (this step is needed because the DAQ install code isn't quite right at the time of this writing.
  • do make install-daq-tpt
  • run starttpt to restart the tpt computer.

Enjoy.
  205   Thu Dec 20 02:04:09 2007 AndreyUpdateComputer Scripts / ProgramsNew overnight measurements in XARM and their results

I ran in the daytime/evening time my program, changing the damping gains in suspension "position" degree of freedom for ETMX and ITMX
in the interval from 1.00 to 3.75 with the step 0.25 (see entry # 201).

Now I am running overnight (from 2AM till 9AM) the program changing the gains in the interval from 1.3 to 3.5 with the step 0.20,
12 X 12 = 144 experimental points. I started so late because I fell asleep after my Wednesday evening dinner, then woke up half an hour ago and hurried to the lab.

BELOW: RESULTS OF MEASUREMENTS WERE ADDED ON THURSDAY EVENING, DEC. 20.

All the meaning of the attachments 1-3, 4-6, 7-9, 10-11 is the same as in previous ELOG entries # 195, # 199, # 202, see in those entries which graph corresponds to which coordinate axes orientation.
Attachment 1: RMS-08Hz-Top-View.png
RMS-08Hz-Top-View.png
Attachment 2: RMS-3Hz-Top-View.png
RMS-3Hz-Top-View.png
Attachment 3: RMS-broadband-Top-View.png
RMS-broadband-Top-View.png
Attachment 4: RMS-08Hz-Side_View.png
RMS-08Hz-Side_View.png
Attachment 5: RMS-3Hz-Side_View.png
RMS-3Hz-Side_View.png
Attachment 6: RMS-broadband-Side_View.png
RMS-broadband-Side_View.png
Attachment 7: RMS-08Hz-Q_I-Q_E-Axes.png
RMS-08Hz-Q_I-Q_E-Axes.png
Attachment 8: RMS-3Hz-Q_I-Q_E-Axes.png
RMS-3Hz-Q_I-Q_E-Axes.png
Attachment 9: RMS-broadband-Q_I-Q_E-Axes.png
RMS-broadband-Q_I-Q_E-Axes.png
Attachment 10: Accelerometer-ETMX.png
Accelerometer-ETMX.png
Attachment 11: Accelerometer-ITMX.png
Accelerometer-ITMX.png
  204   Wed Dec 19 20:28:27 2007 AndreyDAQPEMNames for all 6 accelerometers have been changed

I eventually changed the names for all 6 accelerometers (see my ELOG entry # 172 from Dec. 05 about my intent to do that).

I removed the word "BS" from their names,
and I changed the word combination "ACC_BS_EAST" in the old name for "ACC_ITMX" in the new name;
as well "ACC_BS_WEST" is now replaced by "ACC_ETMX".
(the reasoning behind such a change should become clear from my ELOG entry #172).

New accelerometer names are:
(note: there are no spaces (nowhere!) in the names of accelerometers, but ELOG replaces ": P" written without a space by a strange symbol Tongue)

C1 : PEM - ACC _ ETMX _ X ;
C1 : PEM - ACC _ ETMX _ Y ;
C1 : PEM - ACC _ ETMX _ Z ;
C1 : PEM - ACC _ ITMX _ X ;
C1 : PEM - ACC _ ITMX _ Y ;
C1 : PEM - ACC _ ITMX _ Z .

One can find them in "C1 : PEM - ACC" in Dataviewer.

  203   Wed Dec 19 16:40:12 2007 steveUpdateSAFETYlaser safety glasses measured
I measured the coarse transission at 1064nm of the 40m safety glasses today.

12 pieces of UVEX # LOTG-YAG/CO2 light green, all plastic construction, ADSORBANT

3 pieces of 6KG5, Scott colored filter glass type,

individual prescription glasses: alan, bob, ben, jay and steve

7 pieces of dual waveleght glasses

These glasses showed 0.00mW transmission out of 170mW Crysta Laser 1064
  202   Wed Dec 19 16:07:37 2007 AndreySummaryComputer Scripts / ProgramsResults of overnight measurements Tue/Wed night (entry #198)

As indicated in ELOG entry 198, I was making overnight measurements during last night from Tuesday to Wednesday.

I was changing the suspension damping gain in ETMX and ITMX in "position" degree of freedom between values of 1.00 and 4.50 with the step 0.25.

Results for RMS of peaks (A) at 0.8Hz, (B) at about 3.0Hz and (C) in the range from 0.6Hz to 3.7Hz ("RMS in a broad interval") are given below:

I plotted three results for RMS in the abovementioned three intervals in three different ways:

1) view from the top in the axes (Q_{ITMX}+Q_{ETMX})/2 and (Q_{ITMX}-Q_{ETMX}) -> first three graphs (attachments 1 -3);

2) view from the side in the same sum- and difference-axes -> next three graphs (attachments 4-6);

3) view from the side in Q_{ITMX} and Q_{ETMX} axes -> next three graphs (attachments 7-9)

Attachments 10 and 11 show ratios of accelerometer signals at different times of the night/morning.


A little discussion about these graphs:

1) The areas of minima and of rapid growth are the same for all the measurements during all three nights.

2) Tonight there was a strange spike for the values of Q_{ETMX}=2.5 and Q_{ITMX}=4.0. I interpret that as an error of experiment.

3) On all the plots from all three nights there is a wide area of minimum on the plots for RMS at 0.8Hz and for "RMS in the broad interval",
and the graph for "RMS at 3Hz" indicates a clearer minimum in a localized area for Q_{ITMX}=2+-1, Q_{ETMX}=2+-1. Note that this area 2+-1
is included into the wide region of minimum for "RMS at 0.8Hz" and "RMS in a broad range".

Therefore, my guess at this stage is that we can choose the optimized value of suspension damping gains for both Q_{ITMX} and Q_{ETMX} somewhere
around 2+-1. I would like to make another overnight measurement (tonight) in that narrowed region with a small step to have more certainty.

By the way, I realized that I was a little bit careless and at some plots Q_I stands for {Q_ITMX}, and Q_E stands for Q_{ETMX}.
Attachment 1: RMS_08Hz_Top_view.png
RMS_08Hz_Top_view.png
Attachment 2: RMS_3Hz_Top_view.png
RMS_3Hz_Top_view.png
Attachment 3: RMS_broad_Top_view.png
RMS_broad_Top_view.png
Attachment 4: RMS_08Hz_Side_view.png
RMS_08Hz_Side_view.png
Attachment 5: RMS_3Hz_Side_view.png
RMS_3Hz_Side_view.png
Attachment 6: RMS_broadband_Side_view.png
RMS_broadband_Side_view.png
Attachment 7: RMS_08Hz_Q_I-Q_E-axes.png
RMS_08Hz_Q_I-Q_E-axes.png
Attachment 8: RMS_3Hz_Q_I-Q_E-axes.png
RMS_3Hz_Q_I-Q_E-axes.png
Attachment 9: RMS_broadband_Q_I-Q_E-axes.png
RMS_broadband_Q_I-Q_E-axes.png
Attachment 10: Accelerom_ETMX.png
Accelerom_ETMX.png
Attachment 11: Accelerom_ITMX.png
Accelerom_ITMX.png
  201   Wed Dec 19 15:51:00 2007 AndreyUpdateComputer Scripts / ProgramsDaytime measurements in XARM and their results

I was making measurements in XARM for three different nights. All the results agree with each other (I will put the results from the last night soon).

Steve Vass recommended to me to compare those results with the daytime data, in order to see if there is a real necessity to run the scripts overnight or if daytime results will yield similar results.

XARM has been locked, and I am taking measurements today from 3.30PM till 11.30PM.

I will be changing the suspension damping gains in ETMX and ITMX "position" degrees of freedom in the interval from 1.0 to 3.75 with the step 0.25.

BELOW: RESULTS OF MEASUREMENTS WERE ADDED ON THURSDAY, DEC. 20.

All the meaning of the attachments 1-3, 4-6, 7-9, 10-11 is the same as in previous ELOG entries # 195, # 199, # 202, see in those entries which graph corresponds to which coordinate axes orientation.
Attachment 1: RMS-08Hz-Top_View.png
RMS-08Hz-Top_View.png
Attachment 2: RMS-3Hz-Top_View.png
RMS-3Hz-Top_View.png
Attachment 3: RMS-broadband-Top_View.png
RMS-broadband-Top_View.png
Attachment 4: RMS-08Hz-Side-View.png
RMS-08Hz-Side-View.png
Attachment 5: RMS-3Hz-Side_View.png
RMS-3Hz-Side_View.png
Attachment 6: RMS-broadband-Side_View.png
RMS-broadband-Side_View.png
Attachment 7: RMS-08Hz-Q_I-Q_E-Axes.png
RMS-08Hz-Q_I-Q_E-Axes.png
Attachment 8: RMS-3Hz-Side_View.png
RMS-3Hz-Side_View.png
Attachment 9: RMS-broadband-Side_View.png
RMS-broadband-Side_View.png
Attachment 10: Accelerometer_ETMX.png
Accelerometer_ETMX.png
Attachment 11: Accelerometer_ITMX.png
Accelerometer_ITMX.png
  200   Wed Dec 19 11:31:01 2007 steveOmnistructurePEMaircond filter maintenance
Jeff is working on all air condiontion units of the 40m lab
This we do every six months.
Attachment 1: acfilters6m.jpg
acfilters6m.jpg
  199   Tue Dec 18 23:41:00 2007 AndreySummaryComputer Scripts / ProgramsResults of Mon/Tue overnight measurements (entry #194)

Here I inform our community about the results of the measurements of RMS values in XARM during the previous night from Monday to Tuesday (I announced those measurements in ELOG entry #194).

All the plots in today's report seem to agree well with the analogous plots from the night from Saturday to Sunday (those results are given in ELOG entry # 195).

All the intervals in which RMS have been calculated are the same as in yesterday's ELOG entry #195.

I plotted three results for RMS in the abovementioned three intervals in three different ways:

1) view from the top in the axes (Q_{ITMX}+Q_{ETMX})/2 and (Q_{ITMX}-Q_{ETMX}) -> first three graphs (attachments 1 -3);

2) view from the side in the same sum- and difference-axes -> next three graphs (attachments 4-6);

3) view from the side in Q_{ITMX} and Q_{ETMX} axes -> next three graphs (attachments 7-9, also attch. 12), above accelerometer spectra (attachments 10-11).

Also, I compared the ground noise level by comparing spectra of accelerometer signals at different times during that night. As a reminder, before my disease I installed one accelerometer near ITMX and another accelerometer near ETMX (see entries 161 and 172 in ELOG). The plots of ratios of accelerometer signals at different times (pairs of times that were used: 11PM and 2AM, 11PM and 5AM, 11PM and 8AM) are given below, see attachments 10-11. The program was running from 11PM on Monday till 9AM on Tuesday.

As I explained in the previous ELOG entry # 198, tonight I am taking experimental data in the narrowere interval from 1.00 to 4.50 with a smaller step 0.25.
Attachment 1: RMS_08HZ_Top_View.png
RMS_08HZ_Top_View.png
Attachment 2: RMS_3HZ_Top_View.png
RMS_3HZ_Top_View.png
Attachment 3: RMS_broad_Top_View.png
RMS_broad_Top_View.png
Attachment 4: RMS_08HZ_Side_View.png
RMS_08HZ_Side_View.png
Attachment 5: RMS_3HZ_Side_View.png
RMS_3HZ_Side_View.png
Attachment 6: RMS_broad_Side_View.png
RMS_broad_Side_View.png
Attachment 7: RMS_08HZ_Q_E_Q_I_Axes.png
RMS_08HZ_Q_E_Q_I_Axes.png
Attachment 8: RMS_3HZ_Q_E_Q_I_Axes.png
RMS_3HZ_Q_E_Q_I_Axes.png
Attachment 9: RMS_broad_Q_E_Q_I_Axes.png
RMS_broad_Q_E_Q_I_Axes.png
Attachment 10: Accelerometer_ITMX.png
Accelerometer_ITMX.png
Attachment 11: Accelerometer_ETMX.png
Accelerometer_ETMX.png
Attachment 12: RMS_broad_Q_E_Q_I_Axes.png
RMS_broad_Q_E_Q_I_Axes.png
  198   Tue Dec 18 23:27:36 2007 AndreyConfigurationComputer Scripts / ProgramsNew overnight measurements (this night from Tue to Wed)

I am making overnight measurements in XARM tonight.

This is the third night of measurements in XARM, but tonight I am scanning the narrower region between values of damping gain 1.00 and 4.50 with the smaller step 0.25. (for comparison, during two previous measurements the region was between 1.0 and 7.0 with the step 0.5).

I have relocked the XARM before the start of the measurements.

I started running the program at 9.30PM, and it should collect all the data by 9.00AM wednesday morning.

Below are explanations why I chose these different parameters for the interval and step:

I am going to put the results of previous night measurements into the next ELOG entry, and it we be pretty obvious from those graphs that results in XARM from the two previous (different) nights agree well with each other, and the approximate positions of minima and areas of "big growth" of the surfaces are pretty obvious from those graphs. It is clear that RMS are too big for the values of the damping gain bigger than 4.0, and that minima are somewhere near the values of 2.0. But those graphs were too rough to locate a somewhat precise value for the minima. Therefore, I am studying tonight the interval of gains between 1.00 and 4.50 with a smaller step.

A short note how I estimate time that is necessary to collect the experimental data.

there are 15 experimental points for each ETMX and ITMX suspension gains in the interval between 1.00 and 4.50 with the step 0.25. They are: 1.00, 1.25, 1.50, 1.75, 2.00, ..., 3.75, 4.00, 4.25, 4.50. As I am changing both ETMX and ITMX gains, I have an array of 15*15=225 elements.
It takes 3 minutes for each point to collect the data (I wrote the program that way). Therefore, the total time it takes to run the program is: 225*3=675 minutes, or 675/60=11.25 hours, almost 11 and a half hours.
  197   Tue Dec 18 21:31:31 2007 tobinUpdatePSLISS RIN
My measurements of the ISS RIN via the SR785 and via the DAQ disagree considerably. The spectral shapes are very similar, however, so I expect that a constant factor is creeping in somewhere. Measurements taken at the PD DC monitor points using the SR785 attached. There is a lot of excess noise in the 300 Hz - 1 kHz region.
Attachment 1: iss-rin.pdf
iss-rin.pdf
ELOG V3.1.3-